OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle. Find the width of the rectangle. Type a numerical answer in the space provided. Do not include labels or spaces in your answer. wouldnt the answer be 4.5 is the width

OpenStudy (anonymous):

length = w+7 34=2w+2(w+7) (w is the width) 34=4w+14 (distribute 2 into w+7 and add to 2w) 20=4w (subtract 14 from each side) 5=w (divide each side by 4) width = 5, length =12 (using first equation l=w+7)

OpenStudy (anonymous):

tell me if u dont get something?

OpenStudy (anonymous):

but its a 4 in my problem not 7

OpenStudy (anonymous):

yea i know i gave u example do for 4 yourself ?

OpenStudy (anonymous):

can u do it ?

OpenStudy (anonymous):

ha no im a 6th grader taking advanced algebra i dont understand anythhing

OpenStudy (anonymous):

ohk i ll explain

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first length = width+4 inches

OpenStudy (anonymous):

perimeter = length + width

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

2width + 2(width+4inches) length

OpenStudy (anonymous):

just sub in length = w+4

OpenStudy (anonymous):

did u get it?

OpenStudy (anonymous):

its 6.5 right

OpenStudy (anonymous):

yes goo work !

OpenStudy (anonymous):

good*

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

nw:)

OpenStudy (anonymous):

i got one more if you can help

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A 20% acid solution is mixed with a 70% acid solution to get 50 liters of a 40% solution. How much of the 70% solution is used? Type a numerical answer in the space provided. Do not include labels or spaces in your answer.

OpenStudy (anonymous):

can u do it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we will do this question as a+b=c

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x+y=50 liters

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

therfore \[20x+70y=40\times50\]

OpenStudy (anonymous):

did u get what i did there